# The Aging Multiverse: Generating Condition-Aware Facial Aging Tree via Training-Free Diffusion

### v2

@@ Abstract
We introduce the Aging Multiverse, a framework for generating multiple plausible facial aging trajectories from a single image, each conditioned on external factors such as environment, health, and lifestyle. Unlike prior methods that model aging as a single deterministic path, our approach creates an aging tree that visualizes diverse futures. To enable this, we propose a training-free diffusion-based method that balances identity preservation, age accuracy, and condition control. Our key contributions include attention mixing to modulate editing strength and a Simulated Aging Regularization strategy to stabilize edits. Extensive experiments and user studies demonstrate state-of-the-art performance across identity preservation, aging realism, and conditional alignment, outperforming existing editing and age-progression models, which often fail to account for one or more of the editing criteria. By transforming aging into a multi-dimensional, controllable, and interpretable process, our approach opens up new creative and practical avenues in digital storytelling, health education, and personalized visualization.
@@ 3.3 Attention Mixing For Inversion-Editability Trade-off
Here, is a scaling factor that controls the inversion-editing trade-off and serves as an attention alignment matrix. This update ensures that the attention computation reflects both the editing goal and the identity constraint, improving consistency in the final output. Finally, attention layers during editing using DiT is computed using .
@@ 3.4 Simulated Aging Regularization in Attention Layers
Given a target age , we compute a weighting factor based on the relative position between the target age , the upper reference bound (e.g., 70), and the lower reference bound (e.g., 30), and apply this to modulate the inversion features:
@@ 4.2 Evaluation on Celebrity Dataset
Among the flow-based editing baselines, our method achieves the best overall balance between editability and identity preservation. It ranks highest in CLIP-T (0.326), indicating strong alignment with the target prompt, and highest in Age MAE (9.5), demonstrating accurate age transformation. While its ID Similarity score (0.49) is lower than that of RF-Solver-Edit (0.57) and FireFlow (0.51), this is expected: those methods prioritize identity preservation by directly replacing attention features, as discussed in Sec. 3.3. As a result, they achieve near-perfect fidelity to the input image but show minimal responsiveness to editing prompts, as reflected in their significantly lower CLIP-T scores and higher Age MAE values. Visual comparisons in Fig. 3 and 9 further illustrate this trade-off: models with high scores tend to overfit to the input appearance, producing outputs with minimal visual change. Our method, by contrast, produces edits that are both prompt-aligned and identity-aware.
@@ 4.5 Ablation Study
Next, we introduce the text embedding masking strategy (Sec. 3) to disentangle identity and prompt information. This further boosts CLIP-T to 0.317 and reduces AgeMAE to 12.5, but again slightly lowers IDsim, as the edits become more visually distinct. However, as illustrated in Fig. 6, this increase in editing power may occasionally lead to instability or unrealistic outputs. To mitigate these issues, we first incorporate the Key modulation technique to further improve the editing quality as shown in Fig. 6. Finally, our Simulated Aging Regularization (SAR) further enhances edit stability while partially improving identity preservation, as shown in Fig. 7. This final version of our method achieves strong semantic edits with improved robustness, validating the effectiveness of our design choices.

### v3

@@ Abstract
We introduce the Aging Multiverse, a framework for generating multiple plausible facial aging trajectories from a single image, each conditioned on external factors such as environment, health, and lifestyle. Unlike prior methods that model aging as a single deterministic path, our approach creates an aging tree that visualizes diverse futures. To enable this, we propose a training-free diffusion-based method that balances identity preservation, age accuracy, and condition control. Our key contributions include attention mixing to modulate editing strength and a Simulated Aging Regularization strategy to stabilize edits. Extensive experiments and user studies demonstrate state-of-the-art performance across identity preservation, aging realism, and conditional alignment, outperforming existing editing and age-progression models, which often fail to account for one or more of the editing criteria. By transforming aging into a multi-dimensional, controllable, and interpretable process, our approach opens up new creative and practical avenues in digital storytelling, health education, and personalized visualization. Additional visual examples are available on our project website: https://agingmultiverse.github.io/
![[Uncaptioned image]](x2.png)
@@ 3.3 Attention Mixing For Inversion-Editability Trade-off
Here, is a scaling factor that controls the inversion-editing trade-off and serves as an attention alignment matrix. We used in this paper, which we empirically found to strike a good balance between editing and identity preservation. This update ensures that the attention computation reflects both the editing goal and the identity constraint, improving consistency in the final output. Finally, attention layers during editing using DiT is computed using .
@@ 3.4 Simulated Aging Regularization in Attention Layers
Given the age of the input image and a target age , we compute a weighting factor based on the relative position between the input age , the target age , the upper reference bound (e.g., 70), and the lower reference bound (e.g., 30), and apply this to modulate the inversion features:
@@ 4.2 Evaluation on Celebrity Dataset
Among the flow-based editing baselines, our method achieves the best overall balance between editability and identity preservation. It ranks highest in CLIP-T (0.326), indicating strong alignment with the target prompt, and highest in Age MAE (9.5), demonstrating accurate age transformation. While its ID Similarity score (0.49) is lower than that of RF-Solver-Edit (0.57) and FireFlow (0.51), this is expected: those methods prioritize identity preservation by directly replacing attention features, as discussed in Sec. 3.3. As a result, they achieve near-perfect fidelity to the input image but show minimal responsiveness to editing prompts, as reflected in their significantly lower CLIP-T scores and higher Age MAE values. Visual comparisons in Fig. 3 and 4 further illustrate this trade-off: models with high scores tend to overfit to the input appearance, producing outputs with minimal visual change. Our method, by contrast, produces edits that are both prompt-aligned and identity-aware.
@@ 4.5 Ablation Study
Next, we introduce the text embedding masking strategy (Sec. 3) to disentangle identity and prompt information. This further boosts CLIP-T to 0.317 and reduces AgeMAE to 12.5, but again slightly lowers ID Similarity, as the edits become more visually distinct. However, as illustrated in Fig. 8, this increase in editing power may occasionally lead to instability or unrealistic outputs. To mitigate these issues, we first incorporate the Key modulation technique to further improve the editing quality as shown in Fig. 8. Finally, our Simulated Aging Regularization (SAR) further enhances edit stability while partially improving identity preservation, as shown in Fig. 9. This final version of our method achieves strong semantic edits with improved robustness, validating the effectiveness of our design choices.
@@ 5 Conclusion
Acknowledgments
This research was supported in part by Lenovo Research (Morrisville, NC). We gratefully acknowledge the invaluable support and assistance of the members of the Mobile Technology Innovations Lab. This work was also supported in part by the National Science Foundation under Grant No. 2213335.